# Unc-TTP: A Method for Classifying LLM Uncertainty to Improve In-Context Example Selection

### v1

@@ 1 Introduction
“Real knowledge is to know the extent of one’s ignorance.” —— Confucius
@@
In the rapidly evolving field of artificial intelligence (AI), the quest for generic AI assistants to solve problems with truthful and reliable information remains challenging. Although Large Language Models (LLMs) are capable of fulfilling user requests for various downstream tasks, they struggle with admitting their knowledge boundary (Xu et al. 2024). Additionally, they sometimes fabricate statements that are difficult to discern from falsehoods, posing trustworthiness issues (Maynez et al. 2020; Alkaissi and McFarlane 2023; Ji et al. 2023). These challenges push us to consider two critical questions: (1) How can we determine whether the LLM is certain about its answer? (2) Can we leverage the uncertainty property of the model to improve its performance?
@@
Much effort has been dedicated to understanding when an LLM is certain of its responses. One line of work employs white-box access to the internal model information to provide interpretability for deep neural networks (Gal and Ghahramani 2016; Amini et al. 2020; Kuhn, Gal, and Farquhar 2023). These methods become increasingly complex and resource-intensive as models scale to LLMs (Doshi-Velez and Kim 2017; Krishna et al. 2023), and are unsuitable for closed-source LLMs like GPT-3.5 and GPT-4 without the involvement of a white-box proxy model.
@@
To broadly evaluate different LLMs (both open- and close-sourced), another line of work turns to black-box methods by analyzing the input-output behavior of models to infer their internal reasoning process. For instance, some research has employed prompt engineering methods to assess whether LLMs have truly grasped the essence of knowledge (Wang, Yue, and Sun 2023; Turpin et al. 2023a; Xie et al. 2024; Wei et al. 2024a). These studies suggest that LLMs exhibit a sign of sycophancy, tending to waver between answers when exposed to interference. While this behavior highlights the presence of uncertainty, existing studies have merely observed it without leveraging it as an indicator to measure the LLM’s uncertainty. Further research has explored sampling-based methods to evaluate LLM certainty (Wang et al. 2023; Cole et al. 2023; Manakul, Liusie, and Gales 2023b), which assesses the answer consistency between multiple decoding paths. However, we argue that this approach measures uncertainty too loosely. In this paper, we incorporate label injection into the sampling-based method and classify the uncertainty from the consistency perspective, offering a more fine-grained classification.
@@
Several attempts have been made to exploit the uncertainty (or certainty) properties of LLMs to improve their performance. Xu et al. (2024); Xie et al. (2024) collect data where the model’s performance meets reliability expectations, and then train the model for optimizing reliability. However, these methods rely heavily on the annotated data and demand significant computational resources. Mo et al. (2024) introduce an example selection approach to few-shot information extraction by incorporating both correct and incorrect demonstrations, leveraging self-consistency (Wang et al. 2023) to refine example selection. However, given the loose criteria of sampling-based self-consistency, we believe that a stricter classification of instances could further enhance in-context learning (ICL).
@@
Inspired by the uncertainty sampling strategy in active learning (Lewis and Gale 1994; Lewis and Catlett 1994; Settles 2009), we employ uncertainty to enhance LLMs performance via ICL. Specifically, we utilize the uncertainty categories identified by Unc-TTP to guide the selection of in-context examples. Compared to prior research on example selection, which typically relies on time-consuming retrieval-based methods to select the optimal examples for each test instance (Robertson and Zaragoza 2009; Liu et al. 2021; Rubin, Herzig, and Berant 2022; Zhang, Feng, and Tan 2022), our approach offers a one-example-for-all solution. In our method, example selection is completed in a single round, with the selected examples applicable to all test instances. This approach significantly boosts inference efficiency and delivers notable performance gains compared to the retrieval-based method BM25 (Robertson and Zaragoza 2009), which requires carefully selecting examples for each individual test instance. It also surpasses the random and sampling-based method by a large margin.
@@
  • We conduct a detailed analysis of the instance distribution categorized as “certain” and “uncertain” by both the sampling-based method and Unc-TTP across various LLMs. Our experiments reveal that both methods effectively pinpoint areas of uncertainty within the LLMs, while stronger LLMs exhibit greater certainty. Additionally, when LLMs encounter disagreement with users, they tend to choose the neutral option to align with user expectations.
@@ 2.1 The Manifestations of LLM Uncertainty
While LLMs have performed well in various downstream tasks, their reliability is still questioned, particularly in highly specialized domains  (Alkaissi and McFarlane 2023; Shen et al. 2023). Previous work has identified sycophancy and hallucinations as two main reasons for the unreliability of LLMs, which can be attributed to the manifestations of model uncertainty.
@@
The sycophancy of LLMs is that they frequently wrongly admit mistakes and give biased feedback when questioned by the user  (Sharma et al. 2023). By using LLMs for debates (Wang, Yue, and Sun 2023), introducing misleading elements into prompts (Turpin et al. 2023b, a; Xie et al. 2024; Wei et al. 2024a), or injecting the answer before engaging in Chain-of-Thought (CoT) reasoning (Xu, Qi, and Xu 2024), these findings collectively suggest that sycophancy may be a prevalent trait among LLMs, indicating their inability to maintain a consistent stance.
The patterns of sycophancy have been consistently observed across various LLMs and targeted for reduction in previous works (Sharma et al. 2023; Yang et al. 2023; Wei et al. 2024a, b). To take a step forward, we recognize the wavering behavior behind sycophancy as an indicator of uncertainty within LLMs.
@@
LLMs often generate content that conflicts with sources or lacks factual verification, known as hallucinations (Li et al. 2023a; Huang et al. 2023). Xu et al. (2024) argue that it occurs when LLMs attempt to answer questions beyond their knowledge boundaries. To address this, they train a reward model to enable the LLM to refuse to answer questions about which it is uncertain. Similarly, Xiong et al. (2024) attempt to get LLMs to verbalize their confidence in their answers but reveal a tendency towards overconfidence. We attribute these issues to the LLMs’ inability to express their uncertainty without additional assistance. Nevertheless, we can still unveil their uncertainty through their extrinsic manifestations.
@@ 2.2 ICL and Example Selection Strategy
ICL is a paradigm that allows LLMs to learn tasks given only a few examples as demonstrations without updating model parameters (Dong et al. 2023; Brown et al. 2020).
The most critical factor affecting a model’s ICL performance is the demonstration examples (Kumar and Talukdar 2021; Lu et al. 2022). Significant efforts have been made to select the best in-context examples for LLMs. Liu et al. (2021) retrieve in-context examples based on their semantic similarity to test samples. Nguyen and Wong (2023) suggest selecting examples by calculating their influence. Qin et al. (2023) iteratively select diverse yet strongly correlated examples as demonstrations. Robertson and Zaragoza (2009) employ BM25 to select ICL examples based on their relevance scores. However, these methods select in-context examples for each test instance dynamically, which can be resource-intensive.
Some works utilize examples of what LLMs have done wrong to introduce more information into demonstrations. Mo et al. (2024) leverage both correct and incorrect samples, with incorrect samples selected through self-consistency (Wang et al. 2023), to enhance the diversity of ICL demonstrations. Krishna et al. (2023) select examples through a weak proxy model, and amplify LLM performance by leveraging these difficult examples and their post hoc explanation in ICL. However, these methods either involve sampling-based approaches, which we consider less precise for identifying LLM blind spots or rely on proxy models with white-box access. In contrast, we propose a self-guided, consistency-based method by taking advantage of LLM uncertainty.
@@ 3.1 The Uncertainty Tripartite Testing Paradigm
Under the no-label setting, we instruct the LLM with “Your job is to determine whether the text is A or B”, where “A” and “B” are the placeholder for the actual labels in the dataset. We thereby get the model answer that is derived entirely from the LLM’s own knowledge and reasoning ability, without any external guiding interference.
Under the right-label and wrong-label settings, we provide LLM with either a right or wrong label, intending to lure the model into making the decision we are steering it towards. These settings expose areas where the model is not confident and wavers between answers. However, recent work has shown that LLMs can be easily misled into believing in falsehoods and providing answers that align with the user’s misguided intentions (Wang, Yue, and Sun 2023). Therefore, slavishly guiding the model by concatenating the externally injected correct (or incorrect) label with the task prompt does not necessarily reflect the true uncertainty of the model but rather an inclination toward sycophancy. To diminish this possibility, we prompt the model with “Your job is to determine whether the text is A or B by reference to the given label, which presents the ground truth of the text as A or B. Despite having to refer to the provided labels, you should still have your own thinking and do not change your stance so easily”.
@@ 4.1 Experimental Setup
We perform our experiments with GPT-3.5 and GPT-4  (OpenAI 2024), llama-2-7b-chat (Touvron et al. 2023) and mistral-7b-instruct (Jiang et al. 2023), where all generations are done via greedy decoding.
@@
We select three text classification tasks that involve subjective decision-making for our experiments, according to the subjectivity ranking (Li et al. 2023b). Sarcasm Headlines (SH) (Misra and Arora 2023), a high-quality news headlines sarcasm detection dataset collected from two online news websites. Humor Speech (HS) (Annamoradnejad and Zoghi 2020), a dataset consisting of short texts with both humorous and non-humorous content. Financial Phrasebank (FP) (Malo et al. 2014), a sentiment classification task that divides finance-related sentences as positive, negative, or neutral.
@@
Following Qin et al. (2024), we apply self-consistency (Wang et al. 2023) with 3 decoding paths (temperature 0.7) as sampling-based method. We label instances where all three sampling results are right as CertainR. We additionally label instances where all three samplings are wrong as CertainW, and those who sometimes get one or two right as Uncertain, which aligns with the uncertainty classification of Unc-TTP.
@@ 5.1 Experiment Settings
As for the uncertainty category, we enumerate all possible combinations of outcomes in the three samplings, with at most six for the Unc-TTP method. We choose the one category with the highest accuracy on the validation set as the final in-context examples for testing. It should be noted that our approach is a one-example-for-all method. Once the selection of the uncertainty category is made, the same example is applied to all test instances, eliminating the need to retrieve the most compatible one for each test instance as in previous approaches (Robertson and Zaragoza 2009; Liu et al. 2021; Nguyen, Shaker, and Hüllermeier 2022; Qin et al. 2023).
@@
We evaluate our approach in comparison with the following methods for selecting ICL demonstrations:
  • Random randomly selects examples from the training set as ICL demonstrations.
  • BM25 (Robertson and Zaragoza 2009) ranks the training examples based on their relevance to each test instance. The top-ranked examples are then selected as ICL demonstrations.
  • Similarity  (Liu et al. 2021) selects the top-$k$ semantically similar examples from the training set as demonstrations for each test sample. Following Qin et al. (2023), all samples are encoded by Sentence-BERT (Reimers and Gurevych 2019) to obtain representations for calculating the cosine similarity.
  • SelfCheckGPT (Manakul, Liusie, and Gales 2023a), a sampling-based method that applies in the fact-checking task. It uses a rhetorical question on the corresponding answers after the sampling to verify whether the model supports the answer and scores each instance by the consistency of verification. The higher the score, the more likely the model’s answer is hallucinated; otherwise, it is factual. We use the score computed from the three samplings to guide the in-context example selection.
  • Sampling-based method is inspired by self-consistency  (Wang et al. 2023) and employs $q$ decoding paths, followed by majority voting to identify the instances that the LLM is most certain of. We extend this method by classifying the sampling results into CertainR, CertainW, and Uncertain. The Uncertain includes two categories which are correct once or twice out of $q=3$ samplings.
@@ Appendix C Unc-TTP-Guided Uncertainty Sampling Result on Validation Set
Appendix D Limitations
Secondly, we have only evaluated Unc-TTP on subjective tasks. Given highly subjective tasks often require a deep understanding of contextual subtleties, their labels are sometimes ”biased” themselves depending on the annotator decomposition (Li, Lu, and Yin 2022; Li et al. 2023b). This characteristic is more compatible with Unc-TTP, as it can effectively expose areas where LLMs exhibit uncertainty compared to objective reasoning tasks. For the implementation of Unc-TTP on subjective tasks will be a primary focus in our future work.

### v2

@@ Abstract
Unc-TTP: A Method for Classifying LLM Uncertainty
to Improve In-Context Example Selection
Hsiu-Yuan Huang1,2, Zichen Wu1,2, Yutong Yang1,2, Junzhao Zhang1,3, Yunfang Wu1,2thanks:   Corresponding author. 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University, Beijing, China 3School of Software and Microelectronics, Peking University, Beijing, China {huang.hsiuyuan, zhangjunzhao}@stu.pku.edu.cn, {wuzichen, yytpku, wuyf}@pku.edu.cn,
@@ 1 Introduction
"Real knowledge is to know the extent of one’s ignorance." —— Confucius In the rapidly evolving field of artificial intelligence (AI), the quest for generic AI assistants to solve problems with truthful and reliable information remains challenging. Although Large Language Models (LLMs) are capable of fulfilling user requests for various downstream tasks, they struggle with admitting their knowledge boundary (Xu et al., 2024a). Additionally, they sometimes fabricate statements that are difficult to discern from falsehoods, posing trustworthiness issues (Maynez et al., 2020; Alkaissi and McFarlane, 2023; Ji et al., 2023). These challenges push us to consider two critical questions: (1) How can we determine whether the LLM is certain about its answer? (2) Can we leverage the uncertainty property of the model to improve its performance?
@@
Much effort has been dedicated to understanding when an LLM is certain of its responses. One line of work employs white-box access to the internal model information to provide interpretability for deep neural networks (Gal and Ghahramani, 2016; Amini et al., 2020; Kuhn et al., 2023). These methods become increasingly complex and resource-intensive as models scale to LLMs (Doshi-Velez and Kim, 2017; Krishna et al., 2023), and are unsuitable for closed-source LLMs like GPT-3.5 and GPT-4 without the involvement of a white-box proxy model.
@@
To broadly evaluate different LLMs (both open- and close-sourced), another line of work turns to black-box methods by analyzing the input-output behavior of models to infer their internal reasoning process. For instance, some research has employed prompt engineering methods to assess whether LLMs have truly grasped the essence of knowledge (Wang et al., 2023a; Turpin et al., 2023a; Xie et al., 2024; Wei et al., 2024a). These studies suggest that LLMs exhibit a sign of sycophancy, tending to waver between answers when exposed to interference. While this behavior highlights the presence of uncertainty, existing studies have merely observed it without leveraging it as an indicator to measure the LLM’s uncertainty. Further research has explored sampling-based methods to evaluate LLM certainty Wang et al. (2023b); Cole et al. (2023); Manakul et al. (2023b), which assesses the answer consistency between multiple decoding paths. However, we argue that this approach measures uncertainty too loosely. In this paper, we incorporate label injection into the sampling-based method and classify the uncertainty from the consistency perspective, offering a more fine-grained classification.
@@
Several attempts have been made to exploit the uncertainty (or certainty) properties of LLMs to improve their performance. Xu et al. (2024a); Xie et al. (2024) collect data where the model’s performance meets reliability expectations, and then train the model for optimizing reliability. However, these methods rely heavily on the annotated data and demand significant computational resources. Mo et al. (2024) introduce an example selection approach to few-shot information extraction by incorporating both correct and incorrect demonstrations, leveraging self-consistency Wang et al. (2023b) to refine example selection. However, given the loose criteria of sampling-based self-consistency, we believe that a stricter classification of instances could further enhance in-context learning (ICL).
@@
Inspired by the uncertainty sampling strategy in active learning (Lewis and Gale, 1994; Lewis and Catlett, 1994; Settles, 2009), we employ uncertainty to enhance LLMs performance via ICL. Specifically, we utilize the uncertainty categories identified by Unc-TTP to guide the selection of in-context examples. Compared to prior research on example selection, which typically relies on time-consuming retrieval-based methods to select the optimal examples for each test instance (Robertson and Zaragoza, 2009; Liu et al., 2021; Rubin et al., 2022; Zhang et al., 2022), our approach offers a one-example-for-all solution. In our method, example selection is completed in a single round, with the selected examples applicable to all test instances. This approach significantly boosts inference efficiency and delivers notable performance gains compared to the retrieval-based method BM25 (Robertson and Zaragoza, 2009), which requires carefully selecting examples for each individual test instance. It also surpasses the random and sampling-based method by a large margin.
@@
  • We conduct a detailed analysis of the instance distribution categorized as “certain" and “uncertain" by both the sampling-based method and Unc-TTP across various LLMs. Our experiments reveal that both methods effectively pinpoint areas of uncertainty within the LLMs, while stronger LLMs exhibit greater certainty. Additionally, when LLMs encounter disagreement with users, they tend to choose the neutral option to align with user expectations.
@@ 2.1 The Manifestations of LLM Uncertainty
While LLMs have performed well in various downstream tasks, their reliability is still questioned, particularly in highly specialized domains such as medicine  (Alkaissi and McFarlane, 2023; Shen et al., 2023). Previous work has identified sycophancy and hallucinations as two main reasons for the unreliability of LLMs, which can be attributed to the manifestations of model uncertainty.
@@
The sycophancy of LLMs is that they frequently wrongly admit mistakes and give biased feedback when questioned by the user  (Sharma et al., 2023). By using LLMs for debates (Wang et al., 2023a), introducing misleading elements into prompts (Turpin et al., 2023b, a; Xie et al., 2024; Wei et al., 2024a), or injecting the answer before engaging in Chain-of-Thought (CoT) reasoning (Xu et al., 2024b), these findings collectively suggest that sycophancy may be a prevalent trait among LLMs, indicating their inability to maintain a consistent stance.
The patterns of sycophancy have been consistently observed across various LLMs and targeted for reduction in previous works (Sharma et al., 2023; Yang et al., 2023; Wei et al., 2024a, b). To take a step forward, we recognize the wavering behavior behind sycophancy as an indicator of uncertainty within LLMs.
@@
LLMs often generate content that conflicts with sources or lacks factual verification, known as hallucinations (Li et al., 2023a; Huang et al., 2023). Xu et al. (2024a) argue that it occurs when LLMs attempt to answer questions beyond their knowledge boundaries. To address this, they train a reward model to enable the LLM to refuse to answer questions about which it is uncertain. Similarly, Xiong et al. (2024) attempt to get LLMs to verbalize their confidence in their answers but reveal a tendency towards overconfidence. We attribute these issues to the LLMs’ inability to express their uncertainty without additional assistance. Nevertheless, we can still unveil their uncertainty through their extrinsic manifestations.
@@ 2.2 ICL and Example Selection Strategy
ICL is a paradigm that allows LLMs to learn tasks given only a few examples as demonstrations without updating model parameters (Dong et al., 2023; Brown et al., 2020).
The most critical factor affecting a model’s ICL performance is the demonstration examples (Kumar and Talukdar, 2021; Lu et al., 2022). Significant efforts have been made to select the best in-context examples for LLMs. Liu et al. (2021) retrieve in-context examples based on their semantic similarity to test samples. Nguyen and Wong (2023) suggest selecting examples by calculating their influence. Qin et al. (2023) iteratively select diverse yet strongly correlated examples as demonstrations. Robertson and Zaragoza (2009) employ BM25 to select ICL examples based on their relevance scores. However, these methods select in-context examples for each test instance dynamically, which can be resource-intensive.
Some works utilize examples of what LLMs have done wrong to introduce more information into demonstrations. Mo et al. (2024) leverage both correct and incorrect samples, with incorrect samples selected through self-consistency Wang et al. (2023b), to enhance the diversity of ICL demonstrations. Krishna et al. (2023) select examples through a weak proxy model, and amplify LLM performance by leveraging these difficult examples and their post hoc explanation in ICL. However, these methods either involve sampling-based approaches, which we consider less precise for identifying LLM blind spots or rely on proxy models with white-box access. In contrast, we propose a self-guided, consistency-based method by taking advantage of LLM uncertainty.
@@ 3.1 The Uncertainty Tripartite Testing Paradigm
Under the no-label setting, we instruct the LLM with “Your job is to determine whether the text is A or B", where “A" and “B" are the placeholder for the actual labels in the dataset. We thereby get the model answer that is derived entirely from the LLM’s own knowledge and reasoning ability, without any external guiding interference.
Under the right-label and wrong-label settings, we provide LLM with either a right or wrong label, intending to lure the model into making the decision we are steering it towards. These settings expose areas where the model is not confident and wavers between answers. However, recent work has shown that LLMs can be easily misled into believing in falsehoods and providing answers that align with the user’s misguided intentions (Wang et al., 2023a). Therefore, slavishly guiding the model by concatenating the externally injected correct (or incorrect) label with the task prompt does not necessarily reflect the true uncertainty of the model but rather an inclination toward sycophancy. To diminish this possibility, we prompt the model with “Your job is to determine whether the text is A or B by reference to the given label, which presents the ground truth of the text as A or B. Despite having to refer to the provided labels, you should still have your own thinking and do not change your stance so easily".
@@ 4.1 Experimental Setup
We perform our experiments with GPT-3.5 and GPT-4  (OpenAI, 2024), llama-2-7b-chat (Touvron et al., 2023) and mistral-7b-instruct (Jiang et al., 2023), where all generations are done via greedy decoding.
@@
We select three text classification tasks that involve subjective decision-making for our experiments, according to the subjectivity ranking Li et al. (2023b). Sarcasm Headlines (SH) (Misra and Arora, 2023), a high-quality news headlines sarcasm detection dataset collected from two online news websites. Humor Speech (HS) (Annamoradnejad and Zoghi, 2020), a dataset consisting of short texts with both humorous and non-humorous content. Financial Phrasebank (FP) (Malo et al., 2014), a sentiment classification task that divides finance-related sentences as positive, negative, or neutral.
@@
Following Qin et al. (2024), we apply self-consistency Wang et al. (2023b) with 3 decoding paths (temperature 0.7) as sampling-based method. We label instances where all three sampling results are right as CertainR. We additionally label instances where all three samplings are wrong as CertainW, and those who sometimes get one or two right as Uncertain, which aligns with the uncertainty classification of Unc-TTP.
@@ 5.1 Experiment Settings
As for the uncertainty category, we enumerate all possible combinations of outcomes in the three samplings, with at most six for the Unc-TTP method. We choose the one category with the highest accuracy on the validation set as the final in-context examples for testing. It should be noted that our approach is a one-example-for-all method. Once the selection of the uncertainty category is made, the same example is applied to all test instances, eliminating the need to retrieve the most compatible one for each test instance as in previous approaches Robertson and Zaragoza (2009); Liu et al. (2021); Nguyen et al. (2022); Qin et al. (2023).
@@
We evaluate our approach in comparison with the following methods for selecting ICL demonstrations: Random randomly selects examples from the training set as ICL demonstrations. BM25 Robertson and Zaragoza (2009) ranks the training examples based on their relevance to each test instance. The top-ranked examples are then selected as ICL demonstrations. Similarity  Liu et al. (2021) selects the top-$k$ semantically similar examples from the training set as demonstrations for each test sample. Following Qin et al. (2023), all samples are encoded by Sentence-BERT Reimers and Gurevych (2019) to obtain representations for calculating the cosine similarity.
SelfCheckGPT Manakul et al. (2023a), a sampling-based method that applies in the fact-checking task. It uses a rhetorical question on the corresponding answers after the sampling to verify whether the model supports the answer and scores each instance by the consistency of verification. The higher the score, the more likely the model’s answer is hallucinated; otherwise, it is factual. We use the score computed from the three samplings to guide the in-context example selection.
Sampling-based method is inspired by self-consistency  Wang et al. (2023b) and employs $q$ decoding paths, followed by majority voting to identify the instances that the LLM is most certain of. We extend this method by classifying the sampling results into CertainR, CertainW, and Uncertain. The Uncertain includes two categories which are correct once or twice out of $q=3$ samplings.
@@ 6 Conclusion
Limitations
Secondly, we have only evaluated Unc-TTP on subjective tasks. Given highly subjective tasks often require a deep understanding of contextual subtleties, their labels are sometimes "biased" themselves depending on the annotator decomposition (Li et al., 2022, 2023b). This characteristic is more compatible with Unc-TTP, as it can effectively expose areas where LLMs exhibit uncertainty compared to objective reasoning tasks. For the implementation of Unc-TTP on subjective tasks will be a primary focus in our future work.
Ethics Statement
Task Time (min)
Unc-TTP on Train Set 30
1-shot ICL on Validation Set 100
1-shot ICL on Test Set 10
Table 3: The approximate computation time of the experiments. Unc-TTP on Train Set represents the classification of uncertainty categories using Unc-TTP. 1-shot ICL on Validation Set represents the evaluation of 1-shot examples from all Unc-TTP categories. 1-shot ICL on Test Set represents the application of selected uncertainty examples on the test set.
Computational Budget
We run all our experiments on Intel® Xeon® Silver 4310 CPU and A40 with a batch size of 1 and the max token number of 20. The Table 3 shows the approximate average computation time with the 7B model Llama-2 and Mistral as the LLM, using 1 random seed.
Use of AI Assistants
We have employed ChatGPT as a writing assistant, primarily for polishing the text after the initial composition.